OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (anonymous):

Divide everything in the equation by 3 giving you: \[2x+6=4\] then put the equation equal to 4 & -4 -4=2x+6 4=2x+6 solve for x...you'll have two answers

OpenStudy (anonymous):

-3I2x + 6I = -12 I2x + 6I = -12/-3 -(2x + 6) = 4 & 2x + 6 = 4 -2x - 6 = 4 & 2x = 4-6 -2x = 4 + 6 & 2x = -2 -2x = 10 & x = -2/2 x = 10/-2 = -5 & -1

OpenStudy (anonymous):

x = -5 x= -1